Solve the given initial value problem. y + 12y +44y +48y = 0 y(O)= -7, y(0) = 18, y(0) = - 76 y(x) =

Thank you!
Answer #1

y try + 4y + say so > 0+120 +44D +48) y =0 Auriwwery equation m² +12m² uum+48 20 & comt2) (mt) (m+6)=0 m=2,-4, -6 → The gene